Find the axis of symmetry:
y = x^2 - x + 10
I can not seem to understand this method, I did look back at some other ones, but felt it confused me even more, help please, thank you!!!

Answer by stanbon(75887)

Find the axis of symmetry:
y = x^2 - x + 10
Rewrite as follows:
x^2-x+10 = y
Set it up for completing the square:
x^2-x+? = y-10+?
Complete the square and keep the equation balanced.
x^2-x+(1/2)^2 = y-10+(1/2)^2
Factor the left side and simplify the right to get:
(x-(1/2))^2 = y-(39/4)
The vertex is at (1/2,(39/4))
So the axis of symmetry is x=(1/2)
Cheers,
Stan H.